WORD PROBLEM: Two oil pipes deposit oil into a tank. One of the pipes is three times as large as the other. The larger pipe can fill the tank in 2 hours. How long will it take to fill the tank with both pipes?
I have tried this problem over and over again and usually come up with the answer of 4 which makes no sense because the correct answer should be less than 2.

Answer by nerdybill(7384)

WORD PROBLEM: Two oil pipes deposit oil into a tank. One of the pipes is three times as large as the other. The larger pipe can fill the tank in 2 hours. How long will it take to fill the tank with both pipes?
.
Rate of larger pipe: 1tank/2hours
Rate of smaller pipe: (1/3)(1/2) = 1tank/6hours
.
Let x = time it takes both pipes to fill tank
then
x(1/2 + 1/6) = 1
Multiplying both sides by 6:
x(3 + 1) = 6
4x = 6
x = 6/4
or
x = 1.5 hours